OpenStudy (xapproachesinfinity):

let's if i did this one right prove that \[|\sin a-\sin b |\leq |b-a|\]

OpenStudy (xapproachesinfinity):

so i will let f(x)=sinx which is differentiable at (a,b) so then i can use the MVT

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

yes is a straightforward application of the mvt

OpenStudy (xapproachesinfinity):

some \[a<c<b\] such that \[f(b)-f(a)=f'(c)(b-a)\] abs to it \[|f(a)-f(b)|=|f'(c)||b-a|\] \[|\sin a-\sin b |=|\cos c||b-a|\leq |b-a|\]

OpenStudy (xapproachesinfinity):

hey @misty1212 i guess

OpenStudy (misty1212):

finished

OpenStudy (xapproachesinfinity):

i guess did i forgot something

OpenStudy (xapproachesinfinity):

let me recheck the question hhee

OpenStudy (xapproachesinfinity):

could find the original Q darn it lol

OpenStudy (xapproachesinfinity):

shoud be |a-b| i actually forgot the take care of the sign for that side as well \[|\sin a -\sin b|=\cos c|a-b|\leq|a-b|\] hence proved :)